# Generative Point Cloud Registration

### v1

@@ 3.3 Zero-Shot Texture Consistency Generation
$\begin{split}\tilde{\epsilon}_{\theta}(\mathbf{x}_{t}^{\mathcal{P}};t,\mathbf{c},\mathbf{d}_{\mathcal{P}})\rightarrow\mathbf{x}_{t-1}^{\mathcal{P}},\\ \tilde{\epsilon}_{\theta}(\mathbf{x}_{t}^{\mathcal{Q}};t,\mathbf{c},\mathbf{d}_{\mathcal{Q}})\rightarrow\mathbf{x}_{t-1}^{\mathcal{Q}},\end{split}$ (2)
@@
Based on this motivation, we establish Match-ControlNet, an improved ControlNet variant for matching-specific, conditional image generation. Following Sec. 3.2, we still take depth maps derived by point clouds as conditional images so as to inherit ControlNet’s 2D-3D geometric consistency generation capability. Additionally, we introduce two key designs: coupled conditional denoising and coupled prompt guidance to achieve the cross-view texture consistency generation. The details of these two designs are as follows:
@@
To enable effective cross-view message passing without any finetuning (i.e., zero-shot), we propose an efficient coupled conditional denoising scheme for joint, interactive source and target image generations. Specifically, we expand the noisy latent representation $\mathbf{x}_{t}^{\mathcal{P}(\mathcal{Q})}$ with shape $H^{\prime},W^{\prime},d]$ to a coupled one $\mathbf{x}_{t}^{\mathcal{PQ}}$ with the extended shape $2H^{\prime},W^{\prime},d]$. Also, we vertically concatenate the source and target depth maps into a coupled one $\mathbf{D}_{\mathcal{PQ}}\in\mathbb{R}^{2H\times W\times 1}$, and further emply the ControlNet’s zero convolutions to encode it into the condition features $\mathbf{d}_{\mathcal{PQ}}\in\mathbb{R}^{2H^{\prime}\times W^{\prime}\times d}$. Consequently, without any architectural modifications or parameter fine-tuning, the original conditional denoiser can be directly employed for our coupled conditional denoising:
$\tilde{\epsilon}_{\theta}\left(\mathbf{x}_{t}^{\mathcal{PQ}};t,\mathbf{c},\mathbf{d}_{\mathcal{PQ}}\right)\rightarrow\mathbf{x}_{t-1}^{\mathcal{PQ}},$ (3)
@@
Coupled Prompt Guidance. Although the aforementioned coupled denoising generative mechanism has provided the essential components for texture consistency generation, the denoiser still fails to produce the image pair with expected cross-view texture consistency. The core reason is that the denoiser does not know what kind of image the user expects to generate, and we need to tell it what to do. Our important finding is that when we use a specific prompt (named coupled prompt) as below to tell our Match-ControlNet to produce the vertically stacked images with consistent layout and elements:
Generate two vertically stacked images that are captured from the different viewpoints in a same scene. The images should feature the same environment—whether indoor or outdoor, like a living room, office, street, or natural landscape—with very subtle differences between them. Overall, the layout and key elements remain the same.”,
the denoiser can naturally be guided to recover consistent textures without any model fine-tuning. To the best of our knowledge, we are the first to uncover and utilize this inherent capability of the pre-trained ControlNet for zero-shot pairwise image generation enjoying both 2D-3D geometric consistency and cross-view texture consistency.
3.4 Few-Shot Consistency Fine-tuning
Although our zero-shot Match-ControlNet above has demonstrated promising pairwise consistency generation capabilities, Fig. 4 shows that some corresponding regions may still exhibit geometric or texture inconsistency issues. To mitigate it, we further propose a few-shot finetuning mechanism to improve the consistency generation quality of our Match-ControlNet. It’s noted that we only finetune the learnable encoder copy of Match-ControlNet rather than the all parameters to preserve the powerful generation ability of Stable Diffusion. Specifically, we first collect a set of RGB-depth pairs, $\{((\mathbf{I}_{\mathcal{P}},\mathbf{D}_{\mathcal{P}}),(\mathbf{I}_{\mathcal{Q}},\mathbf{D}_{\mathcal{Q}}))_{j}\}$ ($j$ denotes the sample index), as the training data. For each sample pair, we then concatenate its depth maps and RGB images into coupled ones $\{(\mathbf{I}_{\mathcal{P}\mathcal{Q}},\mathbf{D}_{\mathcal{P}\mathcal{Q}})_{j}\}$. Finally, we use the loss function below to finetune the denoiser:
$\vskip-5.69054pt\mathcal{L}=\mathbb{E}_{\mathbf{x}_{t}^{\mathcal{PQ}},t,\tilde{\mathbf{c}},\mathbf{d}_{\mathcal{PQ}},\epsilon\sim\mathcal{N}(0,1)}\left[\left\|\epsilon-\tilde{\epsilon}_{\theta}(\mathbf{x}^{\mathcal{PQ}}_{t},t,\tilde{\mathbf{c}},\mathbf{d}_{\mathcal{PQ}})\right\|_{2}^{2}\right],$ (5)
where $\mathbf{x}^{\mathcal{PQ}}_{t}$ represents the diffused latent representation of the coupled image $\mathbf{I}_{\mathcal{PQ}}$; $\tilde{\mathbf{c}}$ denotes the token sequence of our coupled prompt. Our experiments show that even with a limited number of samples ($\sim$3K), few-shot finetuning can effectively improves the quality of consistency generation.
@@ 3.4 Few-Shot Consistency Fine-tuning
Figure 4: Compared to the zero-shot Match-ControlNet (top), the finetuned Match-ControlNet can tend to achieve higher 2D-3D geometric consistency and the cross-view texture consistency.

### v2

@@ 3.3.2 Zero-Shot Texture Consistency Generation
$\begin{split}\tilde{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{t}^{\mathcal{P}};t,\mathbf{c},\mathbf{d}_{\mathcal{P}})\rightarrow\mathbf{x}_{t-1}^{\mathcal{P}},\\ \tilde{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{t}^{\mathcal{Q}};t,\mathbf{c},\mathbf{d}_{\mathcal{Q}})\rightarrow\mathbf{x}_{t-1}^{\mathcal{Q}},\end{split}$ (2)
@@
Based on this motivation, we establish DepthMatch-ControlNet, an improved ControlNet variant for matching-specific, conditional perspective-view image generation. Following Sec. 3.3.1, we still take depth maps derived by point clouds as conditional images so as to inherit ControlNet’s 2D-3D geometric consistency generation capability. Additionally, we introduce two key designs: coupled conditional denoising and coupled prompt guidance to achieve the cross-view texture consistency generation. The details of these two designs are as follows:
@@
To enable effective cross-view message passing without any finetuning (i.e., zero-shot), we propose an efficient coupled conditional denoising scheme for joint, interactive source and target image generations. Specifically, we expand the noisy latent representation $\mathbf{x}_{t}^{\mathcal{P}(\mathcal{Q})}$ with shape $H^{\prime},W^{\prime},d]$ to a coupled one $\mathbf{x}_{t}^{\mathcal{PQ}}$ with the extended shape $2H^{\prime},W^{\prime},d]$. Also, we vertically concatenate the source and target depth maps into a coupled one $\mathbf{D}_{\mathcal{PQ}}\in\mathbb{R}^{2H\times W\times 1}$, and further employ the ControlNet’s zero convolutions to encode it into the condition features $\mathbf{d}_{\mathcal{PQ}}\in\mathbb{R}^{2H^{\prime}\times W^{\prime}\times d}$. Consequently, without any architectural modifications or parameter fine-tuning, the original conditional denoiser can be directly employed for our coupled conditional denoising:
$\tilde{\boldsymbol{\epsilon}}_{\theta}\left(\mathbf{x}_{t}^{\mathcal{PQ}};t,\mathbf{c},\mathbf{d}_{\mathcal{PQ}}\right)\rightarrow\mathbf{x}_{t-1}^{\mathcal{PQ}},$ (3)
@@
Coupled Prompt Guidance. Although the aforementioned coupled denoising generative mechanism has provided the essential components for texture consistency generation, the denoiser still fails to produce the image pair with expected cross-view texture consistency. The core reason is that the denoiser does not know what kind of image the user expects to generate, and we need to tell it what to do. Our important finding is that when we use a specific prompt (named coupled prompt) as below to tell our DepthMatch-ControlNet to produce the vertically stacked images with consistent layout and elements: “Generate two vertically stacked images that are captured from the different viewpoints in a same scene. The images should feature the same environment—whether indoor or outdoor, like a living room, office, street, or natural landscape—with very subtle differences between them. Overall, the layout and key elements remain the same.”, the denoiser can naturally be guided to recover consistent textures without any model fine-tuning. To the best of our knowledge, we are the first to uncover and utilize this inherent capability of the pre-trained ControlNet for zero-shot pairwise image generation enjoying both 2D-3D geometric consistency and cross-view texture consistency.
@@
Figure 4: Compared to the zero-shot DepthMatch-ControlNet (top), the finetuned DepthMatch-ControlNet can tend to achieve higher 2D-3D geometric consistency and the cross-view texture consistency.
3.3.3 Few-Shot Consistency Fine-tuning
Although our zero-shot DepthMatch-ControlNet above has demonstrated promising pairwise consistency generation capabilities, Fig. 4 shows that some corresponding regions may still exhibit geometric or texture inconsistency issues. To mitigate it, we further propose a few-shot finetuning mechanism to improve the consistency generation quality of our DepthMatch-ControlNet. It’s noted that we only finetune the learnable encoder copy of DepthMatch-ControlNet rather than the all parameters to preserve the powerful generation ability of Stable Diffusion. Specifically, we first collect a set of RGB-depth pairs, $\{((\mathbf{I}_{\mathcal{P}},\mathbf{D}_{\mathcal{P}}),(\mathbf{I}_{\mathcal{Q}},\mathbf{D}_{\mathcal{Q}}))_{j}\}$ ($j$ denotes the sample index), as the training data. For each sample pair, we then concatenate its depth maps and RGB images into coupled ones $\{(\mathbf{I}_{\mathcal{P}\mathcal{Q}},\mathbf{D}_{\mathcal{P}\mathcal{Q}})_{j}\}$. Finally, we use the loss function below to finetune the denoiser:
$\mathcal{L}=\mathbb{E}_{\mathbf{x}_{t}^{\mathcal{PQ}},t,\tilde{\mathbf{c}},\mathbf{d}_{\mathcal{PQ}},\boldsymbol{\epsilon}}\left[\left\|\boldsymbol{\epsilon}-\tilde{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}^{\mathcal{PQ}}_{t},t,\tilde{\mathbf{c}},\mathbf{d}_{\mathcal{PQ}})\right\|_{2}^{2}\right],$ (5)
where the noise variable $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\mathbf{x}^{\mathcal{PQ}}_{t}$ represents the diffused latent representation of the coupled image $\mathbf{I}_{\mathcal{PQ}}$; $\tilde{\mathbf{c}}$ denotes the token sequence of our coupled prompt. Our experiments show that even with a limited number of samples ($\sim$3K), few-shot finetuning can effectively improve the quality of consistency generation.
3.4 LiDARMatch-ControlNet
In addition to generating perspective-view images conditioned on point cloud pairs derived from single-view depth maps in Sec. 3.3, we further propose the LiDARMatch-ControlNet, which takes as input a pair of 360-degree, temporally adjacent LiDAR point clouds and generates paired panoramic RGB images that preserve both geometric and texture consistency for RGB-enhanced LiDAR point cloud registration. Notably, to the best of our knowledge, there are no off-the-shelf models for LiDAR point cloud-to-panoramic image generation. We are the first to successfully realize this task, which requires careful design of various components, including the training data source, 2D/3D data representations, and network architecture, as detailed below.
3.4.1 Panoramic Image Representation
To enable high-quality panoramic image generation from omnidirectional LiDAR point clouds, it is first necessary to determine an appropriate panoramic image representation that effectively captures the complete spatial context of the surrounding scene. Unlike conventional perspective-view images, panoramic representations must preserve continuity across a wide field of view while remaining compatible with image generation models.
Existing panoramic formats generally fall into two categories: (i) Multi-view perspective representations, which divide the 360 scene into discrete, forward-facing subviews (e.g., front, left, right, rear). This representation is widely adopted in autonomous driving systems with multi-camera setups and is advantageous for preserving low-distortion views within each subregion [5]. However, the imposed boundaries between adjacent subviews inherently disrupt the spatial continuity of the scene and segment the global context into disjoint fragments. Consequently, downstream models have to rely on explicit view-stitching strategies to fuse cross-view information, which significantly constrains their ability to capture whole geometric structures and consistent contextual dependencies across the entire scene. (ii) Unified panoramic representations, e.g. equirectangular projections, unfold the entire spherical scene into a continuous 2D image domain. This format encodes the full horizontal (azimuth) field of view ($360^{\circ}$) and vertical (polar) extent ($180^{\circ}$), while preserving spatial continuity and maintaining geometric relationships across the entire scene. Such representations have been widely adopted in $360^{\circ}$ scene understanding tasks (e.g., panoramic semantic segmentation and virtual view synthesis) due to their compatibility with 2D convolutional networks and transformer-based vision backbones.
In our framework, we adopt the equirectangular projection-based unified panoramic representation as the image format, due to its ability to preserve spatial continuity for global context modeling and its seamless architectural compatibility with diffusion-based generative backbones such as Stable Diffusion and ControlNet.
3.4.2 Panoramic LiDAR Point Cloud Representation
Following the determination of the panoramic image format, we construct a corresponding panoramic geometric representation from the 360 LiDAR point clouds to serve as the spatial condition for our Panoramic Match-ControlNet. Specifically, we project the raw LiDAR points into an equirectangular range map that is spatially aligned with the panoramic image domain, thus enabling fine-grained, pixel-wise conditioning for generative guidance.
Let us take the source LiDAR point cloud $\mathcal{P}$ as example (the target point cloud $\mathcal{Q}$ is the same), where each 3D point $\mathbf{p}_{i}=(x_{i},y_{i},z_{i})$ is defined in the LiDAR sensor’s coordinate system. We first transform each point into spherical coordinates $(\theta_{i},\phi_{i})$ as follows:
$\theta_{i}=\arccos({z_{i}}/{\|\mathbf{p}_{i}\|_{2}}),\quad\phi_{i}=\arctan 2(y_{i},x_{i}),$ (6)
where $\theta_{i}$ and $\phi_{i}$ denote the elevation and azimuth angles, respectively. We then adopt the equirectangular projection, which uniformly maps angular coordinates onto a $H\times W$ equirectangular grid, to discretize the spherical domain: $u_{i}=\lfloor{W(\phi_{i}+\pi)}/{2\pi}\rfloor,v_{i}=\lfloor{H\theta_{i}}/{\pi}\rfloor$, with $(u_{i},v_{i})$ indicating the pixel location corresponding to the point $\mathbf{p}_{i}$. The resulting equirectangular range map $\mathbf{D}_{\mathcal{P}}^{\text{equi}}\in\mathbb{R}^{H\times W\times 1}$ records the radial distance $\|\mathbf{p}_{i}\|_{2}$ at each pixel $(u_{i},v_{i})$. Similarly, the range map of the target point cloud is denoted as $\mathbf{D}_{\mathcal{Q}}^{\text{equi}}\in\mathbb{R}^{H\times W\times 1}$. In the presence of multiple points mapping to the same pixel, we retain the nearest point to ensure visibility consistency. This projection effectively flattens the irregular, omnidirectional LiDAR point cloud into an image-like structure aligned with the equirectangular RGB format above, facilitating range map-conditioned panoramic image generation.
3.4.3 Panoramic Coupled Denoising Generation
As performed in Sec. 3.3, our LiDARMatch-ControlNet still follows the coupled conditional denoising mechanism to generate geometrically and texturally consistent panoramic image pairs from paired omnidirectional LiDAR point clouds. Specifically, given the source and target equirectangular range maps derived in Sec. 3.4.2, we vertically concatenate them into an unified conditional map $\mathbf{D}^{\text{equi}}_{\mathcal{PQ}}\in\mathbb{R}^{2H\times W\times 1}$. This map is then encoded via ControlNet’s zero convolutions to produce conditional features $\mathbf{d}^{\text{equi}}_{\mathcal{PQ}}\in\mathbb{R}^{2H^{\prime}\times W^{\prime}\times d}$, which is used to condition the denoising process over the stacked latent panoramic representation $\tilde{\mathbf{x}}_{t}^{\mathcal{PQ}}\in\mathbb{R}^{2H^{\prime}\times W^{\prime}\times d}$. This coupling enables cross-view attention within the UNet, allowing mutual feature interaction across the source and target images without modifying the original model architecture:
$\tilde{\boldsymbol{\epsilon}}_{\theta}\left(\tilde{\mathbf{x}}_{t}^{\mathcal{PQ}};t,\tilde{\mathbf{c}},\mathbf{d}^{\text{equi}}_{\mathcal{PQ}}\right)\rightarrow\tilde{\mathbf{x}}_{t-1}^{\mathcal{PQ}},$ (7)
forming the denoising Markov chain: $\tilde{\mathbf{x}}_{T}^{\mathcal{PQ}}\rightarrow\cdots\tilde{\mathbf{x}}_{1}^{\mathcal{PQ}}\rightarrow\tilde{\mathbf{x}}_{0}^{\mathcal{PQ}}$. The initial coupled latent representation $\tilde{\mathbf{x}}_{T}^{\mathcal{PQ}}$ is sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$. Here, $\tilde{\mathbf{c}}$ represents the token sequence of panoramic coupled prompt: “Generate two vertically stacked surround-view panoramas that are captured from a self-driving vehicle. The panoramas should feature the same urban street environment, like wide roads, vehicles, buildings, and trees—with very subtle differences between them. Overall, the layout and key elements remain the same.”, that further promotes the model to generate vertically stacked panoramic images with consistent scene layout and subtle cross-view differences. However, since there is no off-the-shelf pretrained ControlNet conditioned on range maps for panoramic image generation, zero-shot generation, as achieved in the perspective setting (see Sec. 3.3.1 and Sec. 3.3.2), is not feasible. Instead, effective panoramic generation requires a consistency fine-tuning mechanism, which we detail in the following subsection.
TABLE I: Comparison of LiDAR datasets with respect to camera configurations and the azimuthal and polar field of view of panoramic images.
Datasets Camera FoV Azm. FoV Plr.
KITTI [17] 1 stereo cam $360^{\circ} <$$180^{\circ}$
KITTI-360 [26] 1 stereo + 2 fisheye $360^{\circ}$ $120^{\circ}$
Waymo [39] 5 perspective $360^{\circ} <$$180^{\circ}$
nuScenes [5] 5 perspect. + 1 fisheye $360^{\circ}$ $40^{\circ}$
Dur360BEV [52] 1 spherical cam $\textbf{360}^{\circ}$ $\textbf{180}^{\circ}$
3.4.4 Panoramic Consistency Fine-tuning
To achieve high-quality geometric and texture consistency in our LiDARMatch-ControlNet, we similarly adopt a fine-tuning approach as described in Sec. 3.3.3. However, in the panoramic scenario, the selection of appropriate panoramic data for supervision becomes crucial. We have analyzed several widely-used LiDAR datasets in terms of camera setup and field of view (FoV), as summarized in Table I.
Considering that our goal is to achieve comprehensive consistency across omnidirectional scenes, we select the Dur360BEV dataset [52] for fine-tuning since it uniquely provides genuinely panoramic images captured by a spherical camera system. This dataset offers a complete 360 horizontal FoV (azimuth) and a 180° vertical FoV (polar), ensuring that the generated panoramic pairs are supervised by complete and coherent spatial information without any missing views or segmentation, which is crucial for maintaining global consistency. Specifically, we first collect a set of paired RGB panoramas and corresponding LiDAR-derived range maps, denoted as $\{((\mathbf{I}^{\text{equi}}_{\mathcal{P}},\mathbf{D}^{\text{equi}}_{\mathcal{P}}),(\mathbf{I}^{\text{equi}}_{\mathcal{Q}},\mathbf{D}^{\text{equi}}_{\mathcal{Q}}))_{j}\}$. We vertically concatenate each panorama and its corresponding range map into unified coupled formats $\{(\mathbf{I}^{\rm equi}_{\mathcal{PQ}},\mathbf{D}^{\text{equi}}_{\mathcal{PQ}})_{j}\}$. Finally, we fine-tune only the learnable encoder of LiDARMatch-ControlNet using the following loss function:
$\mathcal{L}=\mathbb{E}_{\tilde{\mathbf{x}}_{t}^{\mathcal{PQ}},t,\tilde{\mathbf{c}},\mathbf{d}^{\text{equi}}_{\mathcal{PQ}},\boldsymbol{\epsilon}}\left[\|\boldsymbol{\epsilon}-\tilde{\boldsymbol{\epsilon}}_{\theta}(\tilde{\mathbf{x}}^{\mathcal{PQ}}_{t},t,\tilde{\mathbf{c}},\mathbf{d}_{\mathcal{PQ}}^{\text{equi}})\|_{2}^{2}\right],$ (8)
where the noise $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\tilde{\mathbf{x}}^{\mathcal{PQ}}_{t}$ is the diffused latent representation of the coupled panoramic images $\mathbf{I}^{\text{equi}}_{\mathcal{PQ}}$. Our experimental results indicate that limited fine-tuning samples (approximately 10K panoramic pairs from Dur360BEV [52]) can achieve consistent panoramic image generation. Please refer to Sec. 4.2 for more details on how these paired training samples are constructed.
3.5 Theoretical Analysis of Coupled Denoising Model
In both the aforementioned DepthMatch-ControlNet and LiDARMatch-ControlNet, we have introduced our proposed coupled conditional denoising mechanism in generating cross-view consistent perspective-view and panoramic images. In this section, we provide a theoretical interpretation of how the coupled denoising mechanism effectively models the joint distribution of cross-view consistent images, thereby enabling high-quality consistent generation.
Specifically, let $(\mathbf{x}^{\mathcal{P}},\mathbf{x}^{\mathcal{Q}},\mathbf{d}_{\mathcal{P}},\mathbf{d}_{\mathcal{Q}})\sim p_{data}$ denote the paired clean cross-view perspective (or panoramic) images and depth (or range) map conditions in the latent space ($p_{data}$ represents the sample distribution). The likelihood over the cross-view image pair can be expressed as $\mathbb{E}_{p_{data}}[p_{\theta}(\mathbf{x}^{\mathcal{P}},\mathbf{x}^{\mathcal{Q}}|\mathbf{d}_{\mathcal{P}},\mathbf{d}_{\mathcal{Q}})]$. By concatenating the cross-view images and condition maps into unified variables $\mathbf{x}^{\mathcal{PQ}}=\operatorname{Cat}([\mathbf{x}^{\mathcal{P}};\mathbf{x}^{\mathcal{Q}}])$ and $\mathbf{d}_{\mathcal{PQ}}=\operatorname{Cat}([\mathbf{d}_{\mathcal{P}};\mathbf{d}_{\mathcal{Q}}])$, we can model the joint distribution of cross-view images via the likelihood over the concatenated representation: $\mathbb{E}_{p_{data}}[p_{\theta}(\mathbf{x}^{\mathcal{PQ}}|\mathbf{d}_{\mathcal{PQ}})]$.
The diffusion process progressively introduces the Gaussian noise into this clean image pair, which gradually converts them into the noise distribution $\mathbf{x}_{T}^{\mathcal{PQ}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, forming a Markov chain $\mathbf{x}^{\mathcal{PQ}}=\mathbf{x}_{0}^{\mathcal{PQ}}\rightarrow\mathbf{x}_{1}^{\mathcal{PQ}}\cdots\rightarrow\mathbf{x}_{T}^{\mathcal{PQ}}$ with $\mathbf{x}_{t}^{\mathcal{PQ}}\sim\mathcal{N}(\mathbf{x}_{t}^{\mathcal{PQ}};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}^{\mathcal{PQ}},(1-\bar{\alpha}_{t})\mathbf{I})$.
Then, the reverse process aims to learn a denoising network, a parameterized normal distribution $p_{\theta}(\mathbf{x}^{\mathcal{PQ}}_{t-1}|\mathbf{x}^{\mathcal{PQ}}_{t},\mathbf{d}_{\mathcal{PQ}}):=\mathcal{N}(\mathbf{x}^{\mathcal{PQ}}_{t-1};\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t}^{\mathcal{PQ}},t,\mathbf{d}_{\mathcal{PQ}}),\beta_{t}\mathbf{I})$ (prompt term is removed for simplicity), to progressively denoise the noisy data $\mathbf{x}_{T}^{\mathcal{PQ}}$ into the clean one $\mathbf{x}_{0}^{\mathcal{PQ}}$, forming a reverse Markov chain $\mathbf{x}_{T}^{\mathcal{PQ}}\rightarrow\mathbf{x}_{T-1}^{\mathcal{PQ}}\rightarrow\cdots\rightarrow\mathbf{x}_{0}^{\mathcal{PQ}}$. Here, $\boldsymbol{\mu}_{\theta}(\mathbf{x}^{\mathcal{PQ}}_{t},t,\mathbf{d}_{\mathcal{PQ}})$ indicates the parameterized mean of the normal distribution. Following [19], the evidence lower bound of the log likelihood over the training data can be derived as the optimization objective for training the denoising network: $\mathcal{L}_{\text{ELBO}}=$
$\small\begin{split}&\mathbb{E}_{q}\Big[{\log p_{\theta}(\mathbf{x}^{\mathcal{PQ}}_{0}\mid\mathbf{x}^{\mathcal{PQ}}_{1},\mathbf{d}_{\mathcal{PQ}})}-\operatorname{D}_{\rm KL}(q(\mathbf{x}_{T}^{\mathcal{PQ}}|\mathbf{x}_{0}^{\mathcal{PQ}})||p(\mathbf{x}_{T}^{\mathcal{PQ}}))\\ &-\sum_{t>1}{\operatorname{D_{KL}}(q(\mathbf{x}^{\mathcal{PQ}}_{t-1}\mid\mathbf{x}^{\mathcal{PQ}}_{t},\mathbf{x}^{\mathcal{PQ}}_{0})||p_{\theta}(\mathbf{x}^{\mathcal{PQ}}_{t-1}\mid\mathbf{x}^{\mathcal{PQ}}_{t},\mathbf{d}_{\mathcal{PQ}}))}.\end{split}$ (9)
Here, the third term is the core loss function, referred to as the denoising matching loss. Through the reparameterization trick $\mathbf{x}_{0}^{\mathcal{PQ}}=\frac{\mathbf{x}_{t}^{\mathcal{PQ}}-\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}}{\sqrt{\bar{\alpha}_{t}}}$ $\left(\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\right)$, the mean of the posterior distribution $q(\mathbf{x}_{t-1}^{\mathcal{PQ}}|\mathbf{x}_{t}^{\mathcal{PQ}},\mathbf{x}_{0}^{\mathcal{PQ}})$ can be rewritten as $\boldsymbol{\mu}_{q}(\mathbf{x}_{t}^{\mathcal{PQ}},\mathbf{x}_{0}^{\mathcal{PQ}})=\frac{1}{\sqrt{\alpha_{t}}}\mathbf{x}_{t}^{\mathcal{PQ}}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}\sqrt{\alpha_{t}}}\boldsymbol{\epsilon}$. Thus, the mean of the denoising network (i.e., prior distribution) can be modeled as $\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t}^{\mathcal{PQ}},t,\mathbf{d}_{\mathcal{PQ}}):=$
$\small\begin{split}\frac{1}{\sqrt{\alpha_{t}}}\mathbf{x}_{t}^{\mathcal{PQ}}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}\sqrt{\alpha_{t}}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t}^{\mathcal{PQ}},t,\mathbf{d}_{\mathcal{PQ}}).\end{split}$ (10)
Consequently, minimizing the KL divergence between the posterior distribution and the prior distribution in the denoising matching loss can be converted to optimizing the parameters of the noise network $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t}^{\mathcal{PQ}},t,\mathbf{d}_{\mathcal{PQ}})$ to approximate the noise $\boldsymbol{\epsilon}$. The final loss function can be formulated as:
$\mathcal{L}=\mathbb{E}_{\mathbf{x}_{t}^{\mathcal{PQ}},t,\mathbf{d}_{\mathcal{PQ}},\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t}^{\mathcal{PQ}},t,\mathbf{d}_{\mathcal{PQ}})\right\|_{2}^{2}\right],$ (11)
which is equivalent to our practical denoising losses as in Eq. 5 and Eq. 8 (the prompt term is omitted for simplicity). This indicates that our coupled denoising mechanism implicitly models the joint model of the cross-view images for consistency generation. In particular, since the cross-view images inherently exhibit consistency, maximizing the likelihood of their joint distribution through the loss function Eq. 11 above can effectively promote our learned coupled denoising network to generate geometrically and texturally consistent perspective-view/panoramic RGB images, thereby improving the matching quality. Please refer to Appendix A for more details.